OpenStudy (anonymous):

ALGEBRA 2 HELP PLEASE Which prize would you choose and why: get $1 million dollars or earn 1 penny on day 1 and have your salary double each day for 30 days (so day 2’s earnings=$0.02, day 3’s earnings=$0.04, day 4’s earnings = $0.08, …) HELP SHOW WORK . thanks

OpenStudy (turingtest):

for the second option the sequence of your money can be written (in cents) as 1,2,4,8,.... this is a geometric series\[S=2^0,2^1,2^2,2^3,\]or in general, on the n-th day\[S_n=2^n\]on the 30th day, what entry in the sequence is that? how much money is that?

OpenStudy (turingtest):

on the n-th day, the entry is\[S_n=2^{n-1}\]sorry, typo*

OpenStudy (anonymous):

I think you should add all those quantities (what you earn every day), so use a formula for the sum of geometric series Sn = 0.02 (1 - 2^31)/(1 - 2), for n = 30.

OpenStudy (anonymous):

Connections. Huh. I chose the penny prize, becasue eventuallyyou would get more than 1million doolars

OpenStudy (anonymous):

@TuringTest I'm still confused, could you help me a little more, ? thanks

OpenStudy (anonymous):

Day 1: $.01 Day 2: $.02 Day 3: $.04 Day 4: $.08 Day 5: $.16 Day 6: $.32 Day 7: $.64 Day 8: $1.28 Day 9: $2.56 Day 10: $5.12 Day 11: $10.24 Day 12: $20.48 Day 13: $40.96 Day 14: $81.92 Day 15: $163.84 Day 16: $327.68 Day 17: $655.36 Day 18: $1,310.72 Day 19: $2,621.44 Day 20: $5,242.88 Day 21: $10,485.76 Day 22: $20,971.52 Day 23: $41,943.04 Day 24: $83,886.08 Day 25: $167,772.16 Day 26: $335,544.32 Day 27: $671,088.64 Day 28: $1,342,177.28 Day 29: $2,684,354.56 Day 30: $5,368,709.12

OpenStudy (anonymous):

and @viodhora thanks!

OpenStudy (anonymous):

Does that help

OpenStudy (anonymous):

@abdullahnoorulhaqq thank you but can you show the work? message me!

OpenStudy (anonymous):

@abdullahnoorulhaqq i gave you a medal by the way :)

OpenStudy (anonymous):

@abdullahnoorulhaqq THANK YOU SO MUCH

OpenStudy (anonymous):

Welcome